If S = {a, b, c} with P(a) = 7P(b) = 8P(c), find P(a).
Answer #1

Sum of probabilities in a sample space = 1

P(a) = 7P(b)

P(b) = P(a)/7

P(a) = 8P(c)

P(c) = P(a)/8

Therefore,

P(a) + P(b) + P(c) = 1

P(a) + P(a)/7 + P(a)/8 = 1

56P(a) + 8P(a) + 7P(a) = 56

71P(a) = 56

P(a) = 56/71

P(b) = (56/71)/7 = 8/71

P(c) = (56/71)/8 = 7/71